OpenStudy (anonymous):

At the end of each year a self-employed person deposits $1,200 in a retirement account that earns 8 percent annually. How much will be in the account when the individual retires at the age of 60 if the contributions start when the person is 50 years old? Round your answer to the nearest dollar.

OpenStudy (rogue):

\[A(t) = 1200 * 1.08^t\]The investment is for 10 years, so plug in 10 for time and solve :)

OpenStudy (anonymous):

i cant get the square right.

OpenStudy (rogue):

\[A(10) = 1200 * 1.08^{10} = 1200 * 2.15892499 \approx $ 2590.71\]

OpenStudy (anonymous):

B(10) = D ( 1 + r ) ^t = 1,200 * 1.08^10 = $2591